# arXiv:yymm.nnnn

### v1

@@
Spontaneous symmetry breaking is a cornerstone modern physics, defining a wealth of phenomena in condensed-matter and high-energy physics, and beyond. It requires an infinite number of degrees of freedom, and even then, for continuous symmetries, it only works if the spatial dimension is not too low, following the classic results of Coleman, Hohenberg, Mermin and Wagner. While usually discussed in the context of quantum and statistical field theories, and in particular, effective field theories, there are advantages in addressing the same kind of phenomena on discrete geometric structures rather than conventional manifolds. When the space is discretized into a lattice, a lucid picture of conventional spontaneous symmetry breaking springs up, with the ultraviolet issues of continuum quantum field theory out-of-sight, and the key effect, which is infrared in nature, revealed through elementary harmonic oscillator networks. From there, it is natural to generalize lattices to other graphs/networks. In this setting, the presence of spontaneous symmetry breaking is controlled by fractional generalizations of resistance distance and the Kirchhoff index, and most broadly by the spectral dimension. Predictably, because of richness of discrete geometric structures in comparison with continuous manifolds, a broader array of geometries emerge where spontaneous breaking of continuous symmetries is blocked by large fluctuations.
@@ 1 Introduction
The above classic considerations of SSB are for continuous spacetimes. We will generalize them to discrete spatial structures. QFTs on discretized spaces have been considered in a number of contexts. The most prominent case is likely the lattice formulation of gauge theories [19] – originally for ordinary vector gauge fields, but later also for higher-tensor fields [20] and topological field theories [21]. The usual case is on regular lattices, but random lattices have also been considered [22], paving the way to more general discrete geometries. Exploration of simplicial quantum gravity [23] has provided motivations for studies of quantum fields on simplicial complexes [24, 25, 26], and on more general networks [27], while recent years have seen considerable interest in Dirac operators within similar setups [28, 29, 30, 31, 32].
In application to SSB, discretization confers a few distinctive advantages, and the purpose of this essay is to explore these advantages. For conventional SSB in continuous spaces, replacing the geometric background with a hypercubic lattice alleviates all ultraviolet issues typical of QFTs, so that there are no divergences and the computations are straightforward and well-defined. This is a common rationale for considering latticized QFTs in general. In out case, it will reduce the question of whether a symmetry is spontaneously broken to an undergraduate-level exercise: computing explicit expectation values in the ground state of a harmonic oscillator network. Pedagogically, this is a very nice way to look at all the standard textbook results in the continuum, as well as their nonrelativistic and higher-derivative generalizations that emerged in recent years [14, 15, 16, 17, 18]. From this point, generalization to arbitrary networks comes in very naturally, since it amounts to replacing lattice Laplacians with graph Laplacians defined in a way that is standard in graph theory and network science. The structure one obtains as a result is richer, however, than what one sees on hypercubic lattices or in continuous spaces, since there is much more room for defining diverse geometries in terms of network connectivities.
@@ 2.1 Shift symmetry and Goldstone bosons
The two fields are thus decoupled at linear order. The field $\eta$ is massive, and its vacuum state is localized around the field configuration $\eta=0$ (this will be visible, in particular, from our subsequent considerations). The action for the field $\phi$ is exactly the same as (1), up to normalization. The key question of SSB is whether the ground state of $\phi$ is localized around some particular field value, in which case the field-shift symmetry is broken, or if it is not, the symmetry is “restored.” Note that the field-shift symmetry of $\phi$ is exactly identical to the original $U(1)$ phase rotation symmetry of $h$, that is, $h\to e^{ia}h$. The field $\phi$ is nothing but the Goldstone boson associated with SSB, while its shift symmetry is a representation of the original continuous symmetry whose SSB one is trying to examine. (One often hears about the absence or presence of Goldstone bosons as such, but what is meant by that is whether the Goldstone field has conventional properties with a well-defined expectation value and particle excitations, as happens in cases with SSB [8].)
@@ 2.3 The ground state and its fluctuations
Before proceeding with generalizations, it is wise to dwell for a moment on the particularly interesting and natural implementation of field shift symmetries in worldvolume theories of extended solitonic objects. Classical topological solutions with energies localized around lower-dimensional submanifolds (solitons, vortex lines, domain walls) are known to translate into dynamical objects in the corresponding quantum theory [37]. The transverse fluctuations of such objects (for example, vibrations of a quantum vortex line) are described by lower-dimensional massless scalar QFTs living on the worldvolume, with the fields corresponding to the local displacement of the solitonic object in the transverse directions. The field shift symmetry literally becomes translational symmetry in the transverse direction. Because there is no SSB in 1d, this implies that quantum vortex lines behave differently from higher-dimensional extended solitons [38, 39, 40]. In particular, transverse momentum is well-defined and conserved, and the final transverse momentum of the vortex line following a scattering event must reflect that. Failure to account for this effect while expanding around classical solutions corresponding to static vortex lines leads to infrared divergences at higher orders. This is analogous to the much better known recoil effect [38, 41, 42, 43, 44] for fully localized point-like solitons. Evidently, in that case, the “worldline” theory describing the transverse fluctuations is quantum mechanics where SSB is impossible as there are finitely many degrees of freedom. That is why transverse momentum is conserved, and the soliton starts moving (recoils) due to the impact of fundamental quanta in what is known as a meson-soliton scattering process.
@@ 2.4 Nonrelativistic QFTs
with the Laplacian eigenvalue density $\rho(\lambda)$. But this expression can never diverge due to contributions near $\lambda=0$. Whatever singularity $\rho(\lambda)$ may have near the origin, this singularity must be integrable due to the probability normalization of $\rho(\lambda)$. One the other hand, $1/\sqrt{\lambda+m^{2}}$ approaches a constant at the origin and cannot make this singularity any stronger, so that the integral is convergent. Unsurprisingly, massive fields are always localized near the value $\phi=0$ in the field space.
@@ 3.1 Quantum fields on general graphs and graph Laplacians
which agrees with the above definition in terms of quadratic forms.
@@
with an arbitrary symmetric function $w$. The particular choice $w(d,d^{\prime})=1/\sqrt{dd^{\prime}}$ corresponds to the so-called normalized graph Laplacian, used instead of the ordinary (combinatorial) graph Laplacian (42) in some applications. We shall mostly focus, however, on the particularly simple generalization of the conventional lattice discretization of continuous field theories provided by (42).
@@ 3.2 Fractional resistance distance, Kirchhoff index and spectral dimension
On a general graph, there is no equivalent of translation symmetry, and the vertices are in general inequivalent. For that reason, $W_{I}$ defined by (23) depends, in general, on the vertex number $I$. One can even imagine that $W_{I}$ would diverge on some sites and not on others as the graph grows, though admittedly, this situations is somewhat exotic. (As a trivial example, consider an $\ell\times\ell$ two-dimensional square lattice with the endpoints of a one-dimensional lattice of length $O(\ell)$ or more attached to two of its sites. The symmetry will be broken on the two-dimensional component, but restored on the long one-dimensional loop as $\ell$ tends to $\infty$.)
@@
just like in (41). One can recognize the parameter $d_{s}$ as nothing but the spectral dimension of the graph, one of the standard measures that generalize the ordinary topological dimension of Euclidean spaces. The spectral dimension originated in condensed matter literature [49, 50, 51, 52, 53] as a quantification of diffusion properties on discrete geometries. It later migrated into a variety of fields, including discrete quantum gravity [54]. Our discussion makes it clear that it is precisely the spectral dimension, or more precisely, the spectral dimension $d_{s}$ defined through the ordinary (combinatorial) Laplacian, that controls — via the condition (57) — the presence of absence of spontaneous symmetry breaking in quantum field theories on graphs/networks defined by (45).
Before proceeding to discuss the properties of spectral dimension on structured graphs, it is wise to review the simplest case of random graphs: the Erdős-Rényi graphs where pairs of vertices are connected independently with a given probability, or the configuration model where the degree distribution is prescribed, but the graph is otherwise random. Simple settings like this do not provide, however, the power law asymptotics (56), and their effective dimension is effectively infinite, ascertaining spontaneous breaking of the field shift symmetry in all cases. The spectral densities of graph Laplacians in such models have been studied in a number of works [55, 56, 57]. The lucid exposition of [56] shows, in particular, that in the absence of vertices of degree 2, there is a spectral gap, and the spectral density $\rho(\lambda)$ is exactly zero below a certain threshold. If there is a finite fraction of vertices of degree 2, the support of $\rho(\lambda)$ extends all the way to $\lambda=0$, but the asympotics in that region is such that $\rho(\lambda)$ decays faster than any power. All of these cases correspond to an infinite spectral dimension.
@@ 3.3 Fractals
Studies of analytic properties of functions on such regular fractals, including the spectra and eigenfunctions of their Laplacians, have attracted considerable attention from mathematicians in recent decades [59, 60]. For the simplest Serpiński gasket built by iterative insertion of triangle motifs, the spectral dimension is well understood and equals $2\ln 3/\ln 5\approx 1.365$ [61, 62], which is comparable to but different from the more familiar Hausdorff dimension $\ln 3/\ln 2\approx 1.585$. This spectral dimension will then determine which field theories possess ground states localized in the field space that break the shift symmetry, and which ones avoid symmetry breaking.
@@ 3.4 Networks
In [70], one finds a prominent discussion of networks with heterogeneous coupling strength, where it is not only specified whether or not two nodes are connected, but also there is an indication of the strength of this connection indicated by the ‘link weights’ $w_{IJ}$. Such ‘link weights’ provide for a natural generalization of (42) given by
@@ Acknowledgments
I thank Ignatios Antoniadis for discussions on spontaneous symmetry breaking in de Sitter, Jarah Evslin for multiple discussions on ground states of low-dimensional QFTs and of Gaussian distributions for onsite field values in the vacuum, and Peter Sollich for answering a few questions about reference [70]. This work is supported by the C2F program at Chulalongkorn University and by NSRF via grant number B41G680029, and additionally by a Priority Research Area DigiWorld grant under the Strategic Programme Excellence Initiative at the Jagiellonian University (Kraków, Poland).

### v2

@@
Spontaneous symmetry breaking is a cornerstone of modern physics, defining a wealth of phenomena in condensed-matter and high-energy physics, and beyond. It requires an infinite number of degrees of freedom, and even then, for continuous symmetries, it only works if the spatial dimension is not too low, following the classic results of Coleman, Hohenberg, Mermin and Wagner. While usually discussed in the context of quantum and statistical field theories, and in particular, effective field theories, there are advantages in addressing the same kind of phenomena on discrete geometric structures rather than conventional manifolds. When the space is discretized into a lattice, a lucid picture of conventional spontaneous symmetry breaking springs up, with the ultraviolet issues of continuum quantum field theory out-of-sight, and the key effect, which is infrared in nature, revealed through elementary harmonic oscillator networks. From there, it is natural to generalize lattices to other graphs/networks. In this setting, the presence of spontaneous symmetry breaking is controlled by fractional generalizations of resistance distance and the Kirchhoff index, and most broadly by the spectral dimension. Predictably, because of the richness of discrete geometric structures in comparison with continuous manifolds, a broader array of geometries emerge where spontaneous breaking of continuous symmetries is blocked by large fluctuations.
@@ 1 Introduction
The above classic considerations of SSB are for continuous spacetimes. We will generalize them to discrete spatial structures. QFTs on discretized spaces have been considered in a number of contexts. The most prominent case is likely the lattice formulation of gauge theories [19] – originally for ordinary vector gauge fields, but later also for higher-tensor fields [20] and topological field theories [21]. The usual case is on regular lattices, but random lattices have also been considered [22], paving the way to more general discrete geometries. Explorations of simplicial quantum gravity [23] have provided motivations for studies of quantum fields on simplicial complexes [24, 25, 26], and on more general networks [27], while recent years have seen considerable interest in Dirac operators within similar setups [28, 29, 30, 31, 32], all the way to interacting fermionic QFTs used as probes of the network geometry [33].
In application to SSB, discretization confers a few distinctive advantages, and the purpose of this essay is to explore these advantages. For conventional SSB in continuous spaces, replacing the geometric background with a hypercubic lattice alleviates all ultraviolet issues typical of QFTs, so that there are no divergences and the computations are straightforward and well-defined. This is a common rationale for considering latticized QFTs in general. In our case, it will reduce the question of whether a symmetry is spontaneously broken to an undergraduate-level exercise: computing explicit expectation values in the ground state of a harmonic oscillator network. Pedagogically, this is a very nice way to look at all the standard textbook results in the continuum, as well as their nonrelativistic and higher-derivative generalizations that emerged in recent years [14, 15, 16, 17, 18]. From this point, generalization to arbitrary networks comes in very naturally, since it amounts to replacing lattice Laplacians with graph Laplacians defined in a way that is standard in graph theory and network science. The structure one obtains as a result is richer, however, than what one sees on hypercubic lattices or in continuous spaces, since there is much more room for defining diverse geometries in terms of network connectivities.
Properties of simple QFTs, including SSB, on discrete geometric structures, including networks, are a natural question from a methodological perspective, and it must be settled with clarity. In applications, such QFTs naturally emerge as a representation for matter fields in discrete approaches to quantum gravity. This most commonly involves simplicial complexes [23], but more general networks have also been considered [34, 35, 36]. Speaking of applications to more practical networks science, in the emerging age of quantum information and quantum communications, multiagent quantum information exchange will lead to communication networks with quantum degrees of freedom on their nodes, and some of such protocols will be naturally described by QFTs. SSB will then serve as a diagnostic of whether a large network can sustain coherent global order (a collective choice among symmetry-related states) or whether quantum fluctuations wash out such order.
@@ 2.1 Shift symmetry and Goldstone bosons
The two fields are thus decoupled at linear order. The field $\eta$ is massive, and its vacuum state is localized around the field configuration $\eta=0$ (this will be visible, in particular, from our subsequent considerations). The action for the field $\phi$ is exactly the same as (1), up to normalization. The key question of SSB is whether the ground state of $\phi$ is localized around some particular field value, in which case the field-shift symmetry is broken, or if it is not, the symmetry is ‘restored.’ Note that the field-shift symmetry of $\phi$ is exactly identical to the original $U(1)$ phase rotation symmetry of $h$, that is, $h\to e^{ia}h$. The field $\phi$ is nothing but the Goldstone boson associated with SSB, while its shift symmetry is a representation of the original continuous symmetry whose SSB one is trying to examine. (One often hears about the absence or presence of Goldstone bosons as such, but what is meant by that is whether the Goldstone field has conventional properties with a well-defined expectation value and particle excitations, as happens in cases with SSB [8].)
@@ 2.3 The ground state and its fluctuations
Before proceeding with generalizations, it is wise to dwell for a moment on the particularly interesting and natural implementation of field shift symmetries in worldvolume theories of extended solitonic objects. Classical topological solutions with energies localized around lower-dimensional submanifolds (solitons, vortex lines, domain walls) are known to translate into dynamical objects in the corresponding quantum theory [41]. The transverse fluctuations of such objects (for example, vibrations of a quantum vortex line) are described by lower-dimensional massless scalar QFTs living on the worldvolume, with the fields corresponding to the local displacement of the solitonic object in the transverse directions. The field shift symmetry literally becomes the translational symmetry in the transverse direction. Because there is no SSB in 1d, this implies that quantum vortex lines behave differently from higher-dimensional extended solitons [42, 43, 44]. In particular, transverse momentum is well-defined and conserved, and the final transverse momentum of the vortex line following a scattering event must reflect that. Failure to account for this effect while expanding around classical solutions corresponding to static vortex lines leads to infrared divergences at higher orders. This is analogous to the much better known recoil effect [42, 45, 46, 47, 48] for fully localized point-like solitons. Evidently, in that case, the ‘worldline’ theory describing the transverse fluctuations is quantum mechanics where SSB is impossible as there are finitely many degrees of freedom. That is why transverse momentum is conserved, and the soliton starts moving (recoils) due to the impact of fundamental quanta in what is known as a meson-soliton scattering process.
@@ 2.4 Nonrelativistic QFTs
with the Laplacian eigenvalue density $\rho(\lambda)$. But this expression can never diverge due to contributions near $\lambda=0$. Whatever singularity $\rho(\lambda)$ may have near the origin, this singularity must be integrable due to the probability normalization of $\rho(\lambda)$. On the other hand, $1/\sqrt{\lambda+m^{2}}$ approaches a constant at the origin and cannot make this singularity any stronger, so that the integral is convergent. Unsurprisingly, massive fields are always localized near the value $\phi=0$ in the field space.
@@ 3.1 Quantum fields on general graphs and graph Laplacians
which agrees with the above definition in terms of quadratic forms. Without loss of generality, we assume the graph to be connected, as different disconnected components would have to be treated completely independently to assess the presence of SSB within each component.
@@
with an arbitrary symmetric function $w$. The particular choice $w(d,d^{\prime})=1/\sqrt{dd^{\prime}}$ corresponds to the so-called normalized graph Laplacian, used instead of the ordinary (combinatorial) graph Laplacian (42) in some applications. We shall mostly focus, however, on the particularly simple generalization of the conventional lattice discretization of continuous field theories provided by (42). Graph Laplacians have found countless applications in network science, including their recent prominent use for network renormalization [50].
@@ 3.2 Fractional resistance distance, Kirchhoff index and spectral dimension
On a general graph, there is no counterpart of translation symmetry, and the vertices are in general inequivalent. For that reason, $W_{I}$ defined by (23) depends, in general, on the vertex number $I$. One can even imagine that $W_{I}$ would diverge on some sites and not on others as the graph grows, though admittedly, this situation is somewhat exotic. (As a trivial example, consider an $\ell\times\ell$ two-dimensional square lattice with the endpoints of a one-dimensional lattice of length $O(\ell)$ or more attached to two of its sites. The symmetry will be broken on the two-dimensional component, but restored on the long one-dimensional loop as $\ell$ tends to $\infty$.)
@@
just like in (41). One can recognize the parameter $d_{s}$ as nothing but the spectral dimension of the graph, one of the standard measures that generalize the ordinary topological dimension of Euclidean spaces. The spectral dimension originated in condensed matter literature [54, 55, 56, 57, 58] as a quantification of diffusion properties on discrete geometries. It later migrated into a variety of fields, including discrete quantum gravity [59]. In contemporary network science, the spectral dimension plays a key role in a few hot topics: dynamical propeties of synchronization within the Kuramoto model [60], studies of the ‘network geometry with flavor’ model [61] and its deterministic counterparts such as the Apollonian networks [62]. Our discussion makes it clear that it is precisely the spectral dimension, or more precisely, the spectral dimension $d_{s}$ defined through the ordinary (combinatorial) Laplacian, that controls — via the condition (57) — the presence of absence of spontaneous symmetry breaking in QFTs on graphs/networks defined by (45). Spectral dimension has previously emerged in considerations of the Mermin-Wagner theorem for graphs [63, 64], closely related to the (zero-temperature) QFT case treated here, though the precise accounting differs, as it does between the classic considerations of Coleman and Mermin-Wagner. Further examples of the role played by the spectral dimension for statistical systems can be seen in [65, 66].
Before proceeding to discuss the properties of spectral dimension on structured graphs, it is wise to review the simplest case of random graphs: the Erdős-Rényi graphs where pairs of vertices are connected independently with a given probability, or the configuration model where the degree distribution is prescribed, but the graph is otherwise random. Simple settings like this do not provide, however, the power law asymptotics (56), and their effective dimension is effectively infinite, ascertaining spontaneous breaking of the field shift symmetry in all cases. (To be more precise, we should be talking about the giant connected component of such graphs, while finite disconnected components trivially cannot support SSB since only a finite number of degrees of freedom can reside on such components.) The spectral densities of graph Laplacians in such models have been studied in a number of works [67, 68, 69]. The lucid exposition of [68] shows, in particular, that in the absence of vertices of degree 2, there is a spectral gap, and the spectral density $\rho(\lambda)$ is exactly zero below a certain threshold. If there is a finite fraction of vertices of degree 2, the support of $\rho(\lambda)$ extends all the way to $\lambda=0$, but the asympotics in that region is such that $\rho(\lambda)$ decays faster than any power. All of these cases correspond to an infinite spectral dimension.
@@ 3.3 Fractals
Studies of analytic properties of functions on such regular fractals, including the spectra and eigenfunctions of their Laplacians, have attracted considerable attention from mathematicians in recent decades [71, 72]. For the simplest Sierpiński gasket built by iterative insertion of triangle motifs, the spectral dimension is well understood and equals $2\ln 3/\ln 5\approx 1.365$ [73, 74], which is comparable to but different from the more familiar Hausdorff dimension $\ln 3/\ln 2\approx 1.585$. This spectral dimension will then determine which field theories possess ground states localized in the field space that break the shift symmetry, and which ones avoid SSB.
@@ 3.4 Networks
In [82], one finds a prominent discussion of networks with heterogeneous coupling strengths, where it is not only specified whether or not two nodes are connected, but there is also a prescription for the strength of this connection indicated by the ‘link weights’ $w_{IJ}$. Such ‘link weights’ provide for a natural generalization of (42) given by
@@
Proceeding to further generalizations, one could make the field $\phi_{I}$ complex and replace $(\phi_{I}-\phi_{J})^{2}$ with $|\phi_{I}-e^{i\theta_{IJ}}\phi_{J}|^{2}$, where $\theta_{IJ}=-\theta_{JI}$. This ushers the discussion into the world of magnetic Laplacians [84], as the phases $\theta_{IJ}$ encode transport properties in an external magnetic field. Lattice versions of such systems commonly emerge in condensed matter physics in the presence of magnetic fields, including the famed ‘Hofstadter butterfly’ [85]. In our context, the most striking aspect of this generalization is that it may lead to all-bands-flat systems [86, 87], where the spectral singularities of the Laplacian start looking zero-dimensional since there is no dispersion, with all the implications that has for SSB, while the spatial geometry is that of a higher-dimensional lattice. Of course, products of such all-bands-flat lattices and regular lattices, or more elaborate structures, could be considered, with the corresponding effects on the magnetic Laplacian spectra.
@@ Acknowledgments
I thank Ignatios Antoniadis for discussions on spontaneous symmetry breaking in de Sitter, Jarah Evslin for multiple discussions on ground states of low-dimensional QFTs and of Gaussian distributions for onsite field values in the vacuum, Peter Sollich for answering a few questions about reference [82], Alessandro Vezzani for bringing to my attention the older works dealing with the Mermin-Wagner theorem on graphs, Sergej Flach for a discussion on the all-bands-flat lattices, and the anonymous referees for helping me with establishing better connections to the contemporary network science literature. This work is supported by the C2F program at Chulalongkorn University and by NSRF via grant number B41G680029, and additionally by a Priority Research Area DigiWorld grant under the Strategic Programme Excellence Initiative at the Jagiellonian University (Kraków, Poland).